# PGT-I: Scaling Spatiotemporal GNNs with Memory-Efficient Distributed Training

### v1

@@ Abstract.
Spatiotemporal graph neural networks (ST-GNNs) are powerful tools for modeling spatial and temporal data dependencies. However, their applications have been limited primarily to small-scale datasets because of memory constraints. While distributed training offers a solution, current frameworks lack support for spatiotemporal models and overlook the properties of spatiotemporal data. Informed by a scaling study on a large-scale workload, we present PyTorch Geometric Temporal Index (PGT-I), an extension to PyTorch Geometric Temporal that integrates distributed data parallel training and two novel strategies: index-batching and distributed-index-batching. Our index techniques exploit spatiotemporal structure to construct snapshots dynamically at runtime, significantly reducing memory overhead, while distributed-index-batching extends this approach by enabling scalable processing across multiple GPUs. Our techniques enable the first-ever training of an ST-GNN on the entire PeMS dataset without graph partitioning, reducing peak memory usage by up to 89% and achieving up to a 13.1x speedup over standard DDP with 128 GPUs.  [^1] To appear in the 2025 International Conference for High Performance Computing, Networking, Storage, and Analysis.
@@ 1. Introduction
  • We validate PGT-I’s efficacy through a scaling study with up to 128 GPUs on ALCF’s Polaris supercomputer   [^3] https://www.alcf.anl.gov/polaris using the PeMS dataset and DCRNN. Our results demonstrate that our techniques scale efficiently, achieving up to an 88.33x speedup with 128 GPUs, while also outperforming standard DDP training by up to 13.10x in terms of workflow runtime and reducing memory usage by up to 89%.
@@ 3. Motivating Case Study
We evaluate training behavior using two implementations of DCRNN. To establish a baseline, we utilize the PyTorch DCRNN implementation introduced in Li et al. (2018). Additionally, we modify the existing PGT-DCRNN implementation to support sequence-to-sequence prediction and batching as described in Li et al. (2018). PGT is a popular  (Mohammadiyeh and Neysiani, 2023; Verdone et al., 2024; Lin et al., 2022) open-source Python library designed to facilitate the development and application of ST-GNNs. PGT models offer speedups over many prior benchmark ST-GNN implementations by leveraging optimized PyG implementations of graph operations, message passing, and parallelism (Fey and Lenssen, 2019). The existing PGT-DCRNN example (Rozemberczki et al., 2021a) predicts only a single time step into the future based on an input sequence. To enable sequence-to-sequence prediction, we incorporate a repeated encoding-decoding mechanism; this approach uses the model’s output from the initial input sequence as input for predicting subsequent time steps into the future, producing an output sequence of equal length to the input sequence.
@@ 3.1. Setup
Given that we observe consistent resource usage across epochs, we limit node-hour costs by profiling training for a single epoch. We record the following epoch-level stats: runtime, train mean absolute error (MAE), and validation MAE. Additionally, we capture system memory usage and GPU memory usage on a per-second basis using psutil and pynvml, respectively. We perform testing on the state-of-the-art Polaris supercomputer using a single compute node. Each compute node is equipped with a 2.8 GHz AMD EPYC Milan 7543P 32-core CPU, 512 GB of DDR4 RAM, and four NVIDIA A100 GPUs. The system is interconnected using HPE Slingshot 11 and employs a Dragonfly topology with adaptive routing. We integrate the standard open-source preprocessing implementation (Li et al., 2018, 2021) into both workflows; and in all cases we use the default training, validation, and test split (Li et al., 2018): 70% train, 10% validation, and 20% test. Whenever possible, training uses the hyperparameters recommended by prior work (Mallick et al., 2020); however, because of GPU memory restrictions in the PyTorch DCRNN implementation, we select a batch size of 32 . Since inference is not the focus of this work, we exclude evaluation with the test set from our experiments.
@@ 3.2. Results
The PeMS-All-LA dataset presents significant memory challenges for both the DCRNN and PGT-DCRNN implementations. As shown in table 2, DCRNN exhibits high memory consumption, with a peak system memory usage of 371.25 GB and peak GPU memory usage of 24.84 GB. Although still suffering from high memory usage, the PGT-DCRNN model demonstrates substantial improvements in memory efficiency (see fig. 2), reducing peak system memory usage to 259.84 GB and peak GPU memory usage to just 1.58 GB. The lower system memory usage in PGT-DCRNN is attributed to standardizing data in place. In contrast, the original DCRNN implementation creates a separate Python object that duplicates the data for standardization, resulting in multiple copies existing in memory during preprocessing. PGT-DCRNN also reduces GPU memory usage by leveraging specialized PGT layers and adopting a streamlined model implementation. Notably, PGT-DCRNN not only minimizes resource demands but also achieves over 15.30x faster runtime relative to the original DCRNN model, highlighting its potential as a foundation for integrating additional scalability techniques.
@@
Despite PGT-DCRNN’s improved performance over DCRNN, neither implementation can scale to the full PeMS dataset because of system memory restrictions. As fig. 2 shows, both DCRNN implementations exceed the system memory limit and crash before beginning training. These results highlight that, unlike the typical challenge of excessive GPU memory usage in standard GNNs (Lee et al., 2024), the primary scalability bottleneck for ST-GNNs lies in restrictive CPU memory consumption.
@@ 4. Design
Section 3.2 demonstrates that even when using the full resources of a Polaris compute node, state-of-the-art tools such as PGT are insufficient for training models on large-scale spatiotemporal datasets. These challenges extend beyond DCRNN to the majority of ST-GNN models (Zhao et al., 2020; Li et al., 2021; Zhu et al., 2021; Roy et al., 2021; He et al., 2023; Liu et al., 2024; Huang et al., 2023; Rozemberczki et al., 2021a; Ye et al., 2022; Zhu et al., 2020) because they similarly utilize sequence-to-sequence data and the standard ST-GNN preprocessing pipeline (Li et al., 2018; Rozemberczki et al., 2021a) that results in significant memory growth. Our case study highlights the need for new tools that explicitly address spatiotemporal data growth and enable distributed, large-scale training. Thus, we propose techniques that decrease memory usage and enable distributed training without utilizing graph partitioning, which can negatively impact accuracy (Mallick et al., 2020; Ji et al., 2024; Hamilton et al., 2018). Given its demonstrated performance, we select PGT as our base and propose an extension – PyTorch Geometric Temporal-Index (PGT-I) – designed to overcome current ST-GNN limitations and fully leverage HPC platforms to accelerate model training. PGT-I is a memory-efficient training framework that is fully integrated into the PGT code base and enables distributed training using Dask and Dask-DDP. Our implementation includes two novel techniques:
@@ 5.2. Benefits of GPU-Index-Batching Relative to Index-Batching
As shown in table 4, GPU-index-batching accelerates data preprocessing and model training over index-batching on the PeMS dataset, reducing runtime by 12.70%. Both implementations complete preprocessing in under 25 seconds, with index-batching requiring 23.88 seconds and GPU-index-batching 17.90 seconds. Since preprocessing accounts for only a small fraction of the total training time, we did not prioritize optimizing it for GPU execution. GPU-index-batching’s speedup is due to the elimination of CPU-to-GPU data transfers during training. As a result of storing the dataset in GPU memory, however, GPU-index-batching uses significantly more GPU memory than does index-batching, requiring 18.60 GB relative to index-batching’s 5.44 GB. This is balanced by reducing CPU memory usage by 60.25%. As shown in fig. 6, index-batching causes a spike in CPU memory usage to approximately 45 GB. This initial spike occurs during preprocessing, specifically dataset standardization. In contrast, GPU-index-batching performs this step on the GPU, avoiding the spike and maintaining a lower overall memory footprint.
@@
Index-batching 332.79 min 45.75 GB 5.43 GB
GPU-index-batching 290.54 min 18.18 GB 18.60 GB
@@ 5.3.1. Scaling
Distributed-index-batching demonstrates favorable scaling relative to its single-GPU implementation, reducing runtime by up to 88.33x with 128 GPUs when including preprocessing and up to 99.81x when considering training time alone. As anticipated, the static preprocessing cost and overhead introduced by DDP (e.g., syncing gradients, AllReduce operations to calculate validation accuracy) hinder linear scaling, especially as overall runtime decreases. Our technique achieves near-linear training scaling with 4, 8, 16, and 32 GPUs but falls short with 64 and 128 GPUs since fixed costs constitute a larger proportion of the total runtime.
@@
PGT-I’s runtime efficiency significantly lowers the barrier to performing training with the PeMS dataset, which was previously intractable even on a state-of-the-art compute node. We note that preprocessing time with distributed-index-batching fluctuates because of the need to perform I/O over a shared parallel file system. This is evident in the abnormally high preprocessing time of approximately 30 seconds with 8 and 32 GPUs in contrast to the 10–15 seconds observed in the majority of index-batching experiments. In follow-up experiments with 4 GPUs, we observed preprocessing times ranging from 11 seconds to 28 seconds, corresponding to measured fluctuations in I/O time rather than data preprocessing. Intuitively, these I/O fluctuations exist regardless of the number of workers. Despite these fluctuations, the preprocessing overhead introduced by index-batching remains relatively low compared with the overall training time, accounting for between 0.25% and 5% (4 and 128 GPUs) of overall runtime for distributed-index-batching experiments.
@@ 5.3.2. Comparison with DDP
Distributed-index-batching significantly outperforms the DDP approach, reducing overall runtime by 2.18x to 13.1x with 4 and 128 GPUs, respectively. This speedup arises from the elimination of inter-node communication and periodic CPU-to-GPU memory transfers, since workers can process data in local memory and utilize a single consolidated CPU-to-GPU memory transfer. As noted in section 5.3.1, because each worker performs preprocessing in local memory, distributed-index-batching’s preprocessing time does not scale with the number of GPUs.
Distributed-index-batching demonstrates significantly better scaling than does DDP. As the number of workers increases, communication overhead limits DDP’s scaling. Furthermore, DDP’s preprocessing time remains relatively stable, reaching a maximum of 305 seconds with 128 workers, despite distributing computation across more workers. The increase in preprocessing time with 128 workers is due to the overhead of distributing data across a larger number of workers, which outweighs the benefits of increased preprocessing parallelization with PeMS. However, DDP maintains a smaller memory footprint than distributed-index-batching when using 32, 64, and 128 GPUs. When using 32 workers, for example, DDP reduces its maximum memory footprint to 53.3 GB, falling below distributed-index-batching’s footprint of 80.45 GB.
@@ 5.3.3. Accuracy with Increasing GPUs
Figure 8 presents the training and validation MAE as the number of GPUs increases. As the number of GPUs used for DDP increases, the optimal training and validation MAE also increases, with a value of 1.84 with 1 GPU and a value of 2.46 with 128 GPUs. This effect has been established in prior literature (Chaubard et al., 2024; You et al., 2017; Goyal et al., 2018); averaging gradients over an increasing number of distributed mini-batches, as well as an increasingly large global batch size, can alter convergence dynamics and may result in diminished learning. This presents a trade-off between final accuracy and computational efficiency. However, the decline in accuracy with more GPUs may be offset by significantly improved runtime. Follow-up testing indicated that the majority of the MAE increase was due to the increased global batch size rather than DDP effects, and techniques such as learning rate scaling (You et al., 2017) reduced the increase in MAE with larger global batch sizes.
@@ 5.4. Scaling to Larger-than-Memory Datasets
Global Shuffling 2.120 2.191 2.323
Local Batch Shuffling 2.103 2.090 2.080
@@
To evaluate the runtime performance of our technique, we utilize the larger PeMS dataset, performing a single epoch of training using 4, 8, 16, 32, 64, and 128 GPUs. We compare with DDP modified to include the changes described above. Notably, while index-batching reduces distributed preprocessing time by an average of 56.86 $\pm$ 30.58 seconds relative to DDP across all configurations, it constitutes a small portion of the overall runtime in a full training workflow. Therefore, we omit further analysis of preprocessing time.
@@ 7. Conclusion
In this work we explored best practices for leveraging modern HPC infrastructure in the ST-GNN training workflow. We introduced index-batching, which significantly reduces the memory footprint of single-GPU training without impacting accuracy or runtime. We further optimized index-batching as GPU-index-batching – a method that performs the workflow entirely in GPU memory and lowers memory usage and runtime by 60.25% and 12.70%, respectively, for single-GPU training using the PeMS dataset. We evaluated distributed-index-batching in a distributed environment with up to 128 GPUs, displaying favorable scaling and outperforming DDP by up to 13.10x. Additionally, we proposed an extension to index-batching that utilizes optimized distributed data management techniques to reduce memory usage and epoch runtime by up to 9.00x and 2.28x, respectively, over DDP. We merge our techniques into the state-of-the-art ST-GNN package [^8] https://github.com/benedekrozemberczki/pytorch_geometric_temporal and publish our source code to promote reproducibility. [^9] https://github.com/uw-mad-dash/PGT_Index

### v2

@@ Abstract.
Spatiotemporal graph neural networks (ST-GNNs) are powerful tools for modeling spatial and temporal data dependencies. However, their applications have been limited primarily to small-scale datasets because of memory constraints. While distributed training offers a solution, current frameworks lack support for spatiotemporal models and overlook the properties of spatiotemporal data. Informed by a scaling study on a large-scale workload, we present PyTorch Geometric Temporal Index (PGT-I), an extension to PyTorch Geometric Temporal that integrates distributed data parallel training and two novel strategies: index-batching and distributed-index-batching. Our index techniques exploit spatiotemporal structure to construct snapshots dynamically at runtime, significantly reducing memory overhead, while distributed-index-batching extends this approach by enabling scalable processing across multiple GPUs. Our techniques enable the first-ever training of an ST-GNN on the entire PeMS dataset without graph partitioning, reducing peak memory usage by up to 89% and achieving up to a 11.78x speedup over standard DDP with 128 GPUs.  [^1] To appear in the 2025 International Conference for High Performance Computing, Networking, Storage, and Analysis.
@@ 1. Introduction
  • We validate PGT-I’s efficacy through a scaling study with up to 128 GPUs on ALCF’s Polaris supercomputer   [^3] https://www.alcf.anl.gov/polaris using the PeMS dataset and DCRNN. Our results demonstrate that our techniques scale efficiently, achieving up to a 115.49x reduction in training time with 128 GPUs, while also outperforming standard DDP training by up to 11.78x in terms of workflow runtime and reducing memory usage by up to 89%.
@@ 3. Motivating Case Study
We evaluate training behavior using two implementations of DCRNN. To establish a baseline, we utilize the PyTorch DCRNN implementation introduced in Li et al. (2018). Additionally, we modify the existing PGT-DCRNN implementation to support batching and stepwise sequence-to-sequence prediction. PGT is a popular  (Mohammadiyeh and Neysiani, 2023; Verdone et al., 2024; Lin et al., 2022) open-source Python library designed to facilitate the development and application of ST-GNNs. PGT models offer speedups over many prior benchmark ST-GNN implementations by leveraging optimized PyG implementations of graph operations, message passing, and parallelism (Fey and Lenssen, 2019). To enable sequence-to-sequence prediction, we extend the model to process the input sequence in a stepwise fashion, feeding it one temporal slice at a time. Specifically, our implementation maintains and updates a hidden state across time steps, producing an output at each step that ultimately forms a prediction sequence of equal length to the input. While PGT-DCRNN implements the diffusion convolution operations described in Li et al. (2018), it is a lightweight variant that uses a single spatiotemporal diffusion convolution layer and does not replicate the full behavior of the original model (e.g., the RNN-based encoder-decoder structure). As such, accuracy may differ between the two implementations. Our case study focuses on runtime and memory consumption, as these are critical factors for scalability, and the proposed optimizations (see section 4.1) are applicable to both the original DCRNN and the simplified PGT-DCRNN.
@@ 3.1. Setup
Given that we observe consistent resource usage across epochs, we limit node-hour costs by profiling training for a single epoch. We record the epoch runtime and capture system memory usage and GPU memory usage on a per-second basis using psutil and pynvml, respectively. We perform testing on the state-of-the-art Polaris supercomputer using a single compute node. Each compute node is equipped with a 2.8 GHz AMD EPYC Milan 7543P 32-core CPU, 512 GB of DDR4 RAM, and four NVIDIA A100 GPUs. The system is interconnected using HPE Slingshot 11 and employs a Dragonfly topology with adaptive routing. We integrate the standard open-source preprocessing implementation (Li et al., 2018, 2021) into both workflows; and in all cases we use the default training, validation, and test split (Li et al., 2018): 70% train, 10% validation, and 20% test. Whenever possible, training uses the hyperparameters recommended by prior work (Mallick et al., 2020); however, because of GPU memory restrictions in the PyTorch DCRNN implementation, we select a batch size of 32. Since inference is not the focus of this work, we exclude evaluation with the test set from our experiments.
@@ 3.2. Results
The PeMS-All-LA dataset presents significant memory challenges for both the DCRNN and PGT-DCRNN implementations. As shown in table 2, DCRNN exhibits high memory consumption, with a peak system memory usage of 371.25 GB and peak GPU memory usage of 24.84 GB. Although still suffering from high memory usage, the PGT-DCRNN model demonstrates substantial lower memory usage (see fig. 2), reducing peak system memory usage to 259.84 GB and peak GPU memory usage to just 1.58 GB. The lower system memory usage in PGT-DCRNN is attributed to standardizing data in place. In contrast, the original DCRNN implementation creates a separate Python object that duplicates the data for standardization, resulting in multiple copies existing in memory during preprocessing. PGT-DCRNN also reduces GPU memory usage and achieves a 15.30x reduction in runtime relative to the original DCRNN model, highlighting its potential as a lightweight foundation for integrating additional scalability techniques.
@@
Despite PGT-DCRNN’s lower resource demands compared to DCRNN, neither implementation can scale to the full PeMS dataset because of system memory restrictions. As fig. 2 shows, both DCRNN implementations exceed the system memory limit and crash before beginning training. These results highlight that, unlike the typical challenge of excessive GPU memory usage in standard GNNs (Lee et al., 2024), the primary scalability bottleneck for ST-GNNs lies in restrictive CPU memory consumption.
@@ 4. Design
Section 3.2 demonstrates that even when using the full resources of a Polaris compute node, state-of-the-art tools such as PGT are insufficient for training models on large-scale spatiotemporal datasets. These challenges extend beyond DCRNN to the majority of ST-GNN models (Zhao et al., 2020; Li et al., 2021; Zhu et al., 2021; Roy et al., 2021; He et al., 2023; Liu et al., 2024; Huang et al., 2023; Rozemberczki et al., 2021a; Ye et al., 2022; Zhu et al., 2020) because they similarly utilize sequence-to-sequence data and the standard ST-GNN preprocessing pipeline (Li et al., 2018; Rozemberczki et al., 2021a) that results in significant memory growth. Our case study highlights the need for new tools that explicitly address spatiotemporal data growth and enable distributed, large-scale training. Thus, we propose techniques that decrease memory usage and enable distributed training without utilizing graph partitioning, which can negatively impact accuracy (Mallick et al., 2020; Ji et al., 2024; Hamilton et al., 2018). Given its performance and flexibility, we select PGT as our base and propose an extension – PyTorch Geometric Temporal-Index (PGT-I) – designed to overcome current ST-GNN limitations and fully leverage HPC platforms to accelerate model training. PGT-I is a memory-efficient training framework that is fully integrated into the PGT code base and enables distributed training using Dask and Dask-DDP. Our implementation includes two novel techniques:
@@ 5.2. Benefits of GPU-Index-Batching Relative to Index-Batching
As shown in table 4, GPU-index-batching accelerates data preprocessing and model training over index-batching on the PeMS dataset, reducing runtime by 12.87%. Both implementations complete preprocessing in under 30 seconds, with index-batching requiring 26.05 seconds and GPU-index-batching 19.05 seconds. Since preprocessing accounts for only a small fraction of the total training time, we did not prioritize optimizing it for GPU execution. GPU-index-batching’s speedup is due to the elimination of CPU-to-GPU data transfers during training. As a result of storing the dataset in GPU memory, however, GPU-index-batching uses significantly more GPU memory than does index-batching, requiring 18.60 GB relative to index-batching’s 5.50 GB. This is balanced by reducing CPU memory usage by 60.30%. As shown in fig. 6, index-batching causes a spike in CPU memory usage to approximately 46 GB. This initial spike occurs during preprocessing, specifically dataset standardization. In contrast, GPU-index-batching performs this step on the GPU, avoiding the spike and maintaining a lower overall memory footprint.
@@
Index-batching 333.58 min 45.84 GB 5.50 GB
GPU-index-batching 290.65 min 18.20 GB 18.60 GB
@@ 5.3.1. Scaling
Distributed-index-batching demonstrates favorable scaling relative to its single-GPU implementation, reducing runtime by up to 79.41x with 128 GPUs when including preprocessing and up to 115.49x when considering training time alone. As anticipated, the static preprocessing cost and overhead introduced by DDP (e.g., syncing gradients, AllReduce operations to calculate validation accuracy) hinder linear scaling, especially as overall runtime decreases. Our technique achieves near-linear training scaling with 4, 8, 16, and 32 GPUs but falls short with 64 and 128 GPUs since fixed costs constitute a larger proportion of the total runtime.
@@
PGT-I’s runtime efficiency significantly lowers the barrier to performing training with the PeMS dataset, which was previously intractable even on a state-of-the-art compute node. We note that preprocessing time with distributed-index-batching fluctuates because of the need to perform I/O over a shared parallel file system and Dask setup time. Thisis evident in the abnormally high preprocessing time of approximately 35-40 seconds with 16 and 32 GPUs in contrast to the 10–20 seconds observed in the majority of index-batching experiments. In follow-up experiments with 4 GPUs, we observed preprocessing times ranging from 11 seconds to 32 seconds, corresponding to measured fluctuations in I/O time rather than data preprocessing. These I/O fluctuations exist regardless of the number of workers. Nevertheless, despite I/O variability, PGT-I maintains fast and scalable training performance, making large-scale spatiotemporal modeling more accessible.
@@ 5.3.2. Comparison with DDP
Distributed-index-batching significantly outperforms the DDP approach, reducing overall runtime by 2.16x to 11.78x with 4 and 128 GPUs, respectively. This speedup arises from the elimination of inter-node communication and periodic CPU-to-GPU memory transfers, since workers can process data in local memory and utilize a single consolidated CPU-to-GPU memory transfer. As noted in section 5.3.1, because each worker performs preprocessing in local memory, distributed-index-batching’s preprocessing time does not scale with the number of GPUs.
Distributed-index-batching demonstrates significantly better scaling than does DDP. As the number of workers increases, communication overhead limits DDP’s scaling. Furthermore, DDP’s preprocessing time remains relatively stable, reaching a maximum of 305 seconds with 128 workers, despite distributing computation across more workers. The increase in preprocessing time with 128 workers is due to the overhead of distributing data across a larger number of workers, which outweighs the benefits of increased preprocessing parallelization with PeMS. However, DDP maintains a smaller memory footprint than distributed-index-batching when using 32, 64, and 128 GPUs. When using 32 workers, for example, DDP reduces its maximum memory footprint to 53.3 GB, falling below distributed-index-batching’s footprint of 90.18 GB.
@@ 5.3.3. Accuracy with Increasing GPUs
Figure 8 presents the training and validation MAE as the number of GPUs increases. As the number of GPUs used for DDP increases, the optimal training and validation MAE also increases, with a value of 1.66 with 1 GPU and a value of 2.23 with 128 GPUs. This effect has been established in prior literature (Chaubard et al., 2024; You et al., 2017; Goyal et al., 2018); averaging gradients over an increasing number of distributed mini-batches, as well as an increasingly large global batch size, can alter convergence dynamics and may result in diminished learning. This presents a trade-off between final accuracy and computational efficiency. However, the decline in accuracy with more GPUs may be offset by significantly improved runtime. Follow-up testing indicated that the majority of the MAE increase was due to the increased global batch size rather than DDP effects, and techniques such as learning rate scaling (You et al., 2017) reduced the increase in MAE with larger global batch sizes.
@@ 5.4. Scaling to Larger-than-Memory Datasets
Global Shuffling 1.932 2.008 2.149
Local Batch Shuffling 1.913 1.868 1.833
@@
To evaluate the runtime performance of our technique, we utilize the larger PeMS dataset, performing a single epoch of training using 4, 8, 16, 32, 64, and 128 GPUs. We compare with DDP modified to include the changes described above. Notably, while index-batching reduces distributed preprocessing time by an average of 56.86 $\pm$ 30.58 seconds relative to DDP across all configurations, preprocessing constitutes a small portion of the overall runtime in a full training workflow. Therefore, we omit further analysis of preprocessing time.
@@ 7. Conclusion
In this work we explored best practices for leveraging modern HPC infrastructure in the ST-GNN training workflow. We introduced index-batching, which significantly reduces the memory footprint of single-GPU training without impacting accuracy or runtime. We further optimized index-batching as GPU-index-batching – a method that performs the workflow entirely in GPU memory and lowers memory usage and runtime by 60.30% and 12.87%, respectively, for single-GPU training using the PeMS dataset. We evaluated distributed-index-batching in a distributed environment with up to 128 GPUs, displaying favorable scaling and outperforming DDP by up to 11.78x. Additionally, we proposed an extension to index-batching that utilizes optimized distributed data management techniques to reduce memory usage and epoch runtime by up to 9.00x and 2.28x, respectively, over DDP. We merge our techniques into the state-of-the-art ST-GNN package [^8] https://github.com/benedekrozemberczki/pytorch_geometric_temporal and publish our source code to promote reproducibility. [^9] https://github.com/uw-mad-dash/PGT_Index